### v1

# Scaling up to Multivariate Rational Function Reconstruction

Abstract
I present an algorithm for the reconstruction of multivariate rational functions from black-box probes. The arguably most important application in high-energy physics is the calculation of multi-loop and multi-leg amplitudes, where rational functions appear as coefficients in the integration-by-parts reduction to basis integrals. I show that for a dense coefficient the algorithm is nearly optimal, in the sense that the number of required probes is close to the number of unknowns.
1 Introduction
The reduction of a large number of scalar Feynman integrals to a smaller set of basis (or master) integrals is an almost universal step in precision calculations in quantum field theories. In many cases, it is also among the most challenging parts of the computation, and has therefore seen lots of attention and development over the years.
The current standard approach is to derive integration-by-parts identities [1] for a set of seed integrals with fixed powers of propagators and irreducible scalar products and solve the resulting system of linear relations via Gauss elimination [2]. The result is a subset of the seed integrals, each expressed in terms of a linear combination of basis integrals. The coefficients are rational functions of kinematic invariants and the space-time dimension.
It is often advantageous to insert numerical values for the dimension and the invariants and solve the system over a finite field [3]. This strategy was initially used to quickly eliminate redundant relations [4]. However, solving the system with sufficiently many different probes, i.e. different numeric values for the variables, it is possible to reconstruct the full result [5, 6]. In this way, one avoids large intermediate expressions and can restrict the reconstruction to those coefficients that are actually needed for the final result, for example a scattering amplitude.
This general strategy has been optimised along several directions. Right from the start, one can attempt to find combinations of relations that lead to better behaved systems or even an explicit recursive solution to the reduction problem [7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18]. Since in a finite-field reduction the very same system of linear equations has to be solved many times, it can be worthwhile to record the solution steps once, optimise the recorded sequence, and replay it to rapidly obtain further probes [19]. Other efforts target the reconstruction itself. The complexity of the rational function coefficients depends on the chosen basis integrals, and a judicious choice results in a factorisation between the dimension and the kinematic invariants [20, 21]. Identifying common factors in the coefficients can further reduce the number of probes required for the reconstruction [22, 23, 24, 25]. Finally, the number of required probes depends on the chosen reconstruction method, and various algorithms with different strengths have been explored [26, 27, 28].
In the following, I present an algorithm for “scaling up” from rational function reconstruction in a single variable to the multivariate case. In section 2, I review reconstruction in a single variable using Thiele interpolation [29]. I then discuss a way to generalise the method to the multivariate case in section 3. The intended application is the reconstruction of coefficients in the reduction to basis integrals. In section 4, I apply the algorithm to complex reduction coefficients in a massive four-loop propagator example and in the two-loop amplitude for diphoton plus jet production [30]. I find that the number of required probes is close to optimal when the fraction of vanishing polynomial coefficients in the numerator and the denominator of the rational function is small.
2 Univariate Rational Function Reconstruction
Excellent introductions into the reconstruction of polynomials and rational functions are given in [6, 26]. Let us briefly review the case of a univariate rational function.
We are given a rational function $f$ in a single variable for which we want to find an explicit form
$f(x)=\frac{p_{n}(x)}{p_{d}(x)},$ (1)
where $p_{n},p_{d}$ are unknown polynomials with no common roots. $f$ is a “black box”, meaning that our only piece of information is an algorithm for computing $f(t)$ for any $t$ in the domain of $f$. One strategy is to construct rational interpolations $f_{N}$ for $N$ probes $(t_{i},f(t_{i}))$ with $i=1,\dots,N$. If $N$ is large enough, we then find $f_{N}=f$ with high probability.
We start with a single probe $(t_{1},f(t_{1}))$ and a constant interpolation
$f_{1}(x)=a_{1}.$ (2)
Requiring $f_{1}(t_{1})=f(t_{1})$ we immediately find $a_{1}=f(t_{1})$. We then add a second probe $(t_{2},f(t_{2}))$. If $f_{1}(t_{2})=f(t_{2})$ we note that we found agreement and continue with the next probe. Otherwise, we introduce the interpolation
$f_{2}(x)=a_{1}+\frac{x-t_{1}}{a_{2}}$ (3)
with $a_{1}=f(t_{1})$ as before and $a_{2}=\frac{t_{2}-t_{1}}{f(t_{2})-a_{1}}$. In general, after $N$ independent probes the interpolation has the form [29]
$f_{N}(x)=a_{1}+\frac{x-t_{1}}{a_{2}+\frac{x-t_{2}}{a_{3}+\frac{x-t_{3}}{\dots+\frac{x-t_{N-1}}{a_{N}}}}}.$ (4)
For adding the next probe $(t_{N+1},f(t_{N+1}))$ we compute
$c_{1}=f(t_{N+1}),\qquad c_{i+1}=\frac{t_{N+1}-t_{i}}{c_{i}-a_{i}},$ (5)
for all $i\leq N$ and construct $f_{N+1}$ with $a_{N+1}=c_{N+1}$ and $a_{1},\dots,a_{N}$ taken from $f_{N}$. If the denominator in equation (5) vanishes, the probe adds no new information. We then check if our present interpolation already agrees for this point, i.e. $f_{N}(t_{N+1})=f(t_{N+1})$, and terminate the reconstruction as soon as some chosen number of probes are predicted correctly. Otherwise we continue with the next probe.
In many cases, the computational cost of the reconstruction is dominated by either the evaluation of the probes or by the divisions in the calculation of the auxiliary constants $c_{j}$ in equation (5). In the latter case, the alternative division-free recursion
$n_{1}={}$ $f(t_{N+1}),$ $n_{i+1}={}$ $t_{N+1}-t_{i}d_{i},$ (6)
$d_{1}={}$ $1,$ $d_{i+1}={}$ $n_{i}-a_{i}d_{i},$ (7)
leading to $a_{N+1}=\frac{n_{N+1}}{d_{N+1}}$ can be much more efficient.
Note that the reconstruction is optimal if two conditions are fulfilled. First, the degrees of the numerator and the denominator in $f$ should be equal or the degree of the denominator should be larger by one. Second, the numerator and denominator polynomials should be perfectly dense, with all coefficients non-zero. In this case the number of required probes is equal to the number of unknown coefficients plus the chosen number of probes used for confirmation. Empirically, the rational functions encountered in integration-by-parts reduction without any kinematic invariants are close to ideal with a typical overhead of about 10% in the number of required probes.
3 Scaling up to Multiple Variables
In general, the rational function to be reconstructed has the form
$f(x_{1},\dots,x_{n})=\frac{\sum C_{p_{1},\dots,p_{n}}x_{1}^{p_{1}}\cdots x_{n}^{p_{n}}}{\sum D_{q_{1},\dots,q_{n}}x_{1}^{q_{1}}\cdots x_{n}^{q_{n}}},$ (8)
where the sums run over all powers $0\leq p_{1}\leq P_{1},\dots,0\leq p_{n}\leq P_{n}$ in the numerator and $0\leq q_{1}\leq Q_{1},\dots,0\leq q_{n}\leq Q_{n}$ in the denominator. The degrees $P_{1},\dots,P_{n}$ and $Q_{1},\dots,Q_{n}$ are a priori unknown. To reconstruct a pair $(P_{i},Q_{i})$ of degrees we can set all other variables to some fixed value, i.e. $x_{j}=t_{j}$ for all $j\neq i$, and perform a univariate rational function reconstruction in the remaining free variable $x_{i}$.
Once we know the degrees, we can in principle determine the unknown coefficients $C,D$ by simply solving a linear system of equations. Knowing the value of $f(t_{1},\dots,t_{n})$, we obtain the linear equation
$\sum C_{p_{1},\dots,p_{n}}t_{1}^{p_{1}}\cdots t_{n}^{p_{n}}=f(t_{1},\dots,t_{n})\sum D_{q_{1},\dots,q_{n}}t_{1}^{q_{1}}\cdots t_{n}^{q_{n}}$ (9)
directly from the definition in equation (8). For $N+1$ coefficients $C_{p_{1},\dots,p_{n}},D_{q_{1},\dots,q_{n}}$, we require $N$ probes to express them in terms of a single coefficient which we can set to an arbitrary non-zero value to fix the overall normalisation. If the numerator and denominator polynomials are dense, the reconstruction is optimal in the sense of needing the lowest possible number of probes. However, assuming constant-time arithmetic for the arguments and coefficients, the time complexity for solving the dense linear system is $\mathcal{O}(N^{3})$ with a space complexity of $\mathcal{O}(N^{2})$. In practice it is usually better to use a method that requires more probes but has better scaling behaviour.
The univariate reconstruction based on Thiele interpolation we discussed in section 2 only requires $\mathcal{O}(N)$ space to store the arguments $t_{1},\dots,t_{N}$ and the coefficients $a_{1},\dots,a_{N}$. To determine these coefficients, one needs to calculate $N(N+1)/2=\mathcal{O}(N^{2})$ auxiliary coefficients (cf. equation (5)), each of which can be computed in constant time. Can we generalise that method to the multivariate case while preserving the superior scaling behaviour?
The main idea is to set all of the variables $x_{1},\dots,x_{n}$ to a single variable $x$, scaled to distinct powers so that we can recover the full dependence on $x_{1},\dots,x_{n}$ after the reconstruction. Concretely, we consider the auxiliary function $g(x)=f(x^{\alpha_{1}},\dots,x^{\alpha_{n}})$ with
$\alpha_{1}=1,\qquad\alpha_{i+1}=[1+\max(P_{i},Q_{i})]\alpha_{i}.$ (10)
We then use univariate reconstruction in $x$ to find an explicit form for $g(x)$. For each term of the form $C_{i}x^{i}$ we can formally interpret the power $i$ as a number in a mixed radix numeral system, where the individual digits correspond to the powers $p_{1},\dots,p_{n}$ of $x_{1},\dots,x_{n}$.
Let us consider a simple example for a black-box function $f(x_{1},x_{2})$. Setting $x_{2}$ to a fixed value $t_{2}$ and using univariate reconstruction in $x_{1}$ we find
$f(x_{1},t_{2})=\frac{C_{0}(t_{2})+C_{1}(t_{2})x_{1}}{D_{0}(t_{2})+D_{2}(t_{2})x_{1}^{2}}.$ (11)
We ignore the coefficients depending on $t_{2}$; our only goal was to learn that the largest power of $x_{1}$ is 2. This tells us to set $\alpha_{2}=2+1$, so we introduce the auxiliary function $g(x)=f(x,x^{3})$. From univariate reconstruction we obtain
$g(x)=\frac{1+x+x^{4}}{1+x^{2}+x^{3}}.$ (12)
The last step is to read off the corresponding powers of the original variables $x_{1},x_{2}$. For the exponents in $g$ we have mixed radix notation $1=0_{?}1_{3},2=0_{?}2_{3},3=1_{?}0_{3},4=1_{?}1_{3}$, where the subscript indicates the numeral base of the corresponding position and the base of the leading digit is irrelevant. This tells us that the original function is
$f(x_{1},x_{2})=\frac{1+x_{1}+x_{1}x_{2}}{1+x_{1}^{2}+x_{2}}.$ (13)
The method described so far can suffer from accidental cancellations between numerator and denominator. For example, for $f(x_{1},x_{2})=\frac{x_{2}}{x_{1}}$ we would obtain the auxillary function $g(x)=x$, which would lead us to believe the original function was $f(x_{1},x_{2})=x_{1}$. To prevent this, we additionally shift the rescaled argument by a randomly chosen number.
Let us summarise the algorithm. Given a rational black-box function $f$ in $n$ variables $x_{1},\dots,x_{n}$
  1. 1.
    For each variable $x_{i}$ with $i<n$, find the largest powers $P_{i}$ and $Q_{i}$ in the numerator and denominator. To do this, set all other variables to randomly chosen values, $x_{j}=t_{j}$ for all $j\neq i$, and use univariate reconstruction in $x_{i}$.
  2. 2.
    Compute the scaling powers $\alpha_{1},\dots,\alpha_{n}$ using equation (10) and choose random shifts $s_{1},\dots,s_{n}$. One of the shifts can be set to zero, e.g. $s_{1}=0$.
  3. 3.
    Use univariate reconstruction to find $g(x)$ from probes $(t_{i},f(t_{i}^{\alpha_{1}}+s_{1},\dots,t_{i}^{\alpha_{n}}+s_{n}))$.
  4. 4.
    For each term $C_{i}x^{i}$ in $g(x)$, recover the powers $p_{1},\dots,p_{n}$ of the original variables $x_{1},\dots,x_{n}$ from the mixed-radix digits of $i$. Then, replace $x^{i}\to(x_{1}-s_{1})^{p_{1}}\cdots(x_{n}-s_{n})^{p_{n}}$.
For the main application we have in mind, namely the reconstruction of coefficients in the reduction to basis integrals, one aims to recover many rational functions from probes with the same arguments. One way to use the same arguments $t_{i}^{\alpha_{1}}+s_{1},\dots,t_{i}^{\alpha_{n}}+s_{n}$ for different functions $f,h,\dots$ is to choose the exponents $\alpha_{1},\dots,\alpha_{n}$ according to the maximum powers of the respective variables in any of the numerators and denominators of $f,h,\dots$. Often, the highest powers of all variables will be determined by a single function, such that the maximum number of required probes remains unaffected. The price to pay is that for the simple functions many vanishing coefficients will be reconstructed, increasing the computing time required for the reconstruction itself. Alternatively, different sets of probe arguments can be used for functions of widely disparate complexity.
4 Application to the Reduction to Basis Integrals
Let us now assess the efficiency of the algorithm presented in section 3 in practical applications. For brevity, we will refer to the new method as “scaling” reconstruction. We compare it to an implementation of an algorithm proposed by Cuyt and Lee [31]. This algorithm is described in detail in [6, 26]. We briefly recall the main steps. First, the variables are rescaled with a common factor $t$ and shifted, leading to $(x_{1},\dots,x_{n})$ = $(ty_{1}+s_{1},\dots,ty_{n}+s_{n})$ with $y_{n}=1$. One then performs a univariate rational reconstruction in $t$. Starting from the highest powers, the coefficients of $t^{i}$ are reconstructed as polynomials in $y_{1},\dots,y_{n}$ and transformed back to the original variables. Here, we are mainly interested in dense rational functions and therefore use nested Newton interpolation for the polynomial reconstruction. We call this “dense homogeneous” reconstruction, adopting terminology from [28].
[…]
4.1 Massive Four-Loop Propagator
Our first benchmark point is a coefficient in the differential equation [34, 35] for a four-loop massive propagator. Setting the mass $m=1$, the propagator is a function of $z=p^{2}$, where $p$ is the external four-momentum. One obtains
[…]
with the ellipsis indicating a linear combination of further basis integrals with less complex coefficients. $q(z,d)$ is a dense rational function in $z$ and the space-time dimension $d$. The numerator degrees in $z$ and $d$ are $P_{z}=P_{d}=81$ and the denominator degrees are $Q_{z}=80$ and $Q_{d}=78$. The number of probes per characteristic required by the dense homogeneous and the scaling algorithm are shown in table 1 together with the number of probes a hypothetical optimal algorithm that can determine one unknown coefficient per probe would use. For a full reconstruction, probes are needed in seven different prime fields, with two probes in an eighth field used for confirmation. The scaling method requires $164$ additional probes in the first prime field to determine the degrees in $z$.
Algorithm Number of Probes
Dense Homogeneous 25 7082570825\,70825 708
Scaling 13 4301343013\,43013 430
Optimal 12 7211272112\,72112 721
Table 1: Number of probes per characteristic required to reconstruct a specific coefficient in the differential equation for a four-loop propagator.
We see that the scaling algorithm performs close to optimal, with an overhead of about $6\%$ additional probes. It requires roughly $48\%$ fewer probes than dense homogeneous reconstruction.
4.2 Diphoton Plus Jet Production at Two Loops
Next, let us consider the two-loop amplitude for diphoton plus jet production, taken from [30]. Specifically, we choose the parity-even contribution with a left-handed quark and a gluon in the initial state, a negative gluon helicity, opposite-sign photon helicities, and no closed fermion loops. Denoting the number of colours by $N_{C}$, we select the largest rational coefficient by Mathematica’s ByteCount in the part proportional to $N_{C}^{0}$. We write this coefficient as a rational function in $x_{23},x_{34},x_{45},x_{51}$, where $x_{ij}=\frac{s_{ij}}{s_{12}}$, and
$s_{12}=(p_{1}+p_{2})^{2},s_{23}=(p_{2}-p_{3})^{2},s_{34}=(p_{3}+p_{4})^{2},s_{45}=(p_{4}+p_{5})^{2},s_{51}=(p_{1}-p_{5})^{2}$
are Mandelstam invariants. In this example, the full coefficient can be reconstructed using a single prime field. The number of required probes is shown in table 2.
Algorithm # Probes
Homogeneous 248 688248688248\,688248 688
Scaling 169 132169132169\,132169 132
Optimal 30 4903049030\,49030 490
Table 2: Number of probes required to reconstruct the coefficient in the reduction of the two-loop diphoton plus jet amplitude.
The scaling algorithm introduced in section 3 performs notably better than the homogeneous one, requiring approximately $32\%$ fewer probes. Still, both algorithms are far from optimal for this scenario and algorithms for sparse reconstruction [26, 27, 36, 37, 38] will likely perform much better.
5 Conclusion
I have presented an algorithm for the reconstruction of dense multivariate rational functions. Multiple variables are mapped onto a single variable, using scaling powers and shifts chosen such that the mapping can be reversed. In this way, the problem is reduced to well-known univariate rational reconstruction.
The algorithm is tested on two examples taken from complex reductions to basis integrals, a massive four-loop propagator and and two-loop five-point amplitude. For the dense rational function encountered in the four-loop problem, the required number of probes exceeds the number of unknown coefficients by only about $6\%$. In the sparse two-loop example, the overhead is much larger and the number of probes exceeds the number of non-vanishing coefficients by a factor of more than $5$. Still, in both cases the scaling algorithm compares favourably to dense homogeneous reconstruction [31]. A promising avenue for future research would be to combine the univariate mapping with sparse rational reconstruction in a single variable [39].
Acknowledgements
I thank P. Marquard for enlightening discussions and helpful comments on the manuscript.

### v2

@@
 In many cases, it is also among the most challenging parts of the computation, and has therefore seen lots of attention and development over the years.
-The current standard approach is to derive integration-by-parts identities [1] for a set of seed integrals with fixed powers of propagators and irreducible scalar products and solve the resulting system of linear relations via Gauss elimination [2].
+The current standard approach is to derive integration-by-parts identities [1, 2] for a set of seed integrals with fixed powers of propagators and irreducible scalar products and solve the resulting system of linear relations via Gauss elimination [3].
 The result is a subset of the seed integrals, each expressed in terms of a linear combination of basis integrals.
@@
 In this way, one avoids large intermediate expressions and can restrict the reconstruction to those coefficients that are actually needed for the final result, for example a scattering amplitude.
-This general strategy has been optimised along several directions.
+Further advantages include ease of parallelisation, lower memory usage, and the possibility to optimise the system further after a computationally cheap pilot run.
+This general strategy has been further developed along several directions.
 Right from the start, one can attempt to find combinations of relations that lead to better behaved systems or even an explicit recursive solution to the reduction problem [8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19].
@@
 Identifying common factors in the coefficients can further reduce the number of probes required for the reconstruction [23, 24, 25, 26].
-Finally, the number of required probes depends on the chosen reconstruction method, and various algorithms with different strengths have been explored [26, 27, 28].
+Finally, the number of required probes depends on the chosen reconstruction method, and various algorithms with different strengths have been explored [27, 28, 29, 30].
 In the following, I present an algorithm for “scaling up” from rational function reconstruction in a single variable to the multivariate case.
@@
 In the latter case, the alternative division-free recursion
-$n_{1}={}$ $f(t_{N+1}),$ $n_{i+1}={}$ $t_{N+1}-t_{i}d_{i},$ (6)
+$n_{1}={}$ $f(t_{N+1}),$ $n_{i+1}={}$ $(t_{N+1}-t_{i})d_{i},$ (6)
 $d_{1}={}$ $1,$ $d_{i+1}={}$ $n_{i}-a_{i}d_{i},$ (7)
@@
 Note that the reconstruction is optimal if two conditions are fulfilled.
-First, the degrees of the numerator and the denominator in $f$ should be equal or the degree of the denominator should be larger by one.
+First, the degrees of the numerator and the denominator in $f$ should be equal or the degree of the numerator should be larger by one.
 Second, the numerator and denominator polynomials should be perfectly dense, with all coefficients non-zero.
@@
 To reconstruct a pair $(P_{i},Q_{i})$ of degrees we can set all other variables to some fixed value, i.e.
-$x_{j}=t_{j}$ for all $j\neq i$, and perform a univariate rational function reconstruction in the remaining free variable $x_{i}$.
+$x_{j}=t_{j}$ for all $j\neq i$, and perform a univariate rational function reconstruction in the remaining free variable $x_{i}$ [7, 27, 33].
 Once we know the degrees, we can in principle determine the unknown coefficients $C,D$ by simply solving a linear system of equations.
@@
 To prevent this, we additionally shift the rescaled argument by a randomly chosen number.
+[^1] This random shift is also used with a slightly different purpose in the reconstruction algorithm by Cuyt and Lee [7, 27, 33].
+There, the goal is to ensure a unique structure and uniform coefficient normalisation of the reconstructed function.
+Spurious cancellations have to involve two or more different variables.
+We therefore expect to avoid them by having at least one shifted variable in each possible combination, i.e.
+we shift each variable except one.
 Let us summarise the algorithm. Given a rational black-box function $f$ in $n$ variables $x_{1},\dots,x_{n}$
@@
 Alternatively, different sets of probe arguments can be used for functions of widely disparate complexity.
+One of the main advantages of numerical reduction is ease of parallelisation.
+This requires that subsequent probes can be chosen without having to wait for the outcome of feeding earlier probes into the reconstruction algorithm.
+In this respect, algorithms for the reconstruction of dense rational functions tend to perform better than methods aimed at sparse rational functions with many vanishing coefficients, as for the latter the probe selection typically has to be adjusted dynamically.
+The presented algorithm mostly decouples the seed choice from the reconstruction progress.
+For $n$ variables, the selection strategy has to be updated $n$ times — after determining the powers of each of the first n1n-1 variables and once more to use the final rescaled arguments.
 4 Application to the Reduction to Basis Integrals
@@
 For brevity, we will refer to the new method as “scaling” reconstruction.
-We compare it to an implementation of an algorithm proposed by Cuyt and Lee [31].
+We compare it to an algorithm proposed by Cuyt and Lee [33].
 This algorithm is described in detail in [7, 27].
@@
 Starting from the highest powers, the coefficients of $t^{i}$ are reconstructed as polynomials in $y_{1},\dots,y_{n}$ and transformed back to the original variables.
-Here, we are mainly interested in dense rational functions and therefore use nested Newton interpolation for the polynomial reconstruction.
-We call this “dense homogeneous” reconstruction, adopting terminology from [28].
-[…]
+For the comparison we use state-of-the-art implementations in the public codes FireFly [27, 29, 34] and FiniteFlow [28, 35].
+The comparison code and the example rational functions in computer-readable form are available from https://github.com/a-maier/scaling-rec.
+For the scaling algorithm, the reconstruction is first performed over a number of prime fields $\mathbb{Z}_{P}$, using arithmetic algorithms taken from NTL [36].
+We start with P=1 152 921 504 606 846 883P=\$1\,152\,921\,504\,606\,846\,883\$ and move to the next smaller prime numbers as needed.
+The resulting coefficients are lifted to a higher characteristic with the Chinese remainder theorem, specifically Bézout’s identity.
+The actual rational coefficients are then reconstructed from the finite-field integers via an algorithm by Wang [37].
+Again, details are given in [7, 27].
+In principle, the reconstruction can be simplified tremendously exploiting the structure of the result [23, 24, 25].
+Both FireFly and FiniteFlow can factorise the numerator and denominator to a certain degree.
+FiniteFlow determines the minimal degree in each variable to automatically factor out common monomials.
+FireFly optionally performs univariate factorisation to identify any factors depending on a single variable.
+Since we are mainly interested in assessing the underlying reconstruction algorithms, we disable FireFly’s factorisation in the following comparisons.
+As there is no option to switch off factorisation in FiniteFlow, we additionally compare the reconstruction after removing all monomial factors from the function to be reconstructed.
 4.1 Massive Four-Loop Propagator
@@
 with the ellipsis indicating a linear combination of further basis integrals with less complex coefficients.
-$q(z,d)$ is a dense rational function in $z$ and the space-time dimension $d$.
-The numerator degrees in $z$ and $d$ are $P_{z}=P_{d}=81$ and the denominator degrees are $Q_{z}=80$ and $Q_{d}=78$.
-The number of probes per characteristic required by the dense homogeneous and the scaling algorithm are shown in table 1 together with the number of probes a hypothetical optimal algorithm that can determine one unknown coefficient per probe would use.
-For a full reconstruction, probes are needed in seven different prime fields, with two probes in an eighth field used for confirmation.
-The scaling method requires $164$ additional probes in the first prime field to determine the degrees in $z$.
-Algorithm Number of Probes
-Dense Homogeneous 25 7082570825\,70825 708
-Scaling 13 4301343013\,43013 430
-Optimal 12 7211272112\,72112 721
-Table 1: Number of probes per characteristic required to reconstruct a specific coefficient in the differential equation for a four-loop propagator.
-We see that the scaling algorithm performs close to optimal, with an overhead of about $6\%$ additional probes. It requires roughly $48\%$ fewer probes than dense homogeneous reconstruction.
+$q(z,d)$ is a rational function in $z$ and the space-time dimension $d$, where the numerator degrees in $z$ and $d$ are $P_{z}=P_{d}=81$ and the denominator degrees are $Q_{z}=80$ and $Q_{d}=78$.
+The numbers of probes required for reconstructing the function over the first characteristic are shown in table 1 for the scaling algorithm and FireFly together with a hypothetical optimal algorithm that can determine one unknown coefficient per probe.
+For a full rational function reconstruction probes are needed in several additional prime fields.
+Since different implementations vary vastly in the amount of reused information we refrain from a quantitative comparison.
+Method Number of Probes
+This work   13 59413\,594
+FireFly   16 37316\,373
+Optimal   12 72112\,721
+Table 1: Number of probes required to reconstruct a specific coefficient in the differential equation for a four-loop propagator over the first characteristic.
+Comparing the “Optimal” entry of table 1 to the total number of 13 12313\,123 monomials in the ansatz given by equation (8) we observe that $q(z,d)$ is dense in the sense that about $97\%$ of the coefficients in the ansatz are non-zero.
+We see that the scaling algorithm performs close to optimal, with an overhead of about $7\%$ additional probes.
+In comparison, FireFly requires approximately $29\%$ more probes.
+The denominator of the reconstructed function contains an overall factor of $d^{1}z^{3}$.
+The improvement gained by identifying and removing this factor is illustrated in table 2, where we now also include FiniteFlow.
+The FiniteFlow entry does not include a few hundred probes used to determine the overall degree of the rational function and the degrees with respect to the individual variables from Thiele interpolation, c.f.
+section 2 [40].
+The other entries count the total number of function evaluations.
+Method Number of Probes
+This work   13 59413\,594
+FireFly   16 02016\,020
+FiniteFlow $\gtrsim{}$18\,205$\$
+Optimal   12 72112\,721
+Table 2: Number of probes required to reconstruct a specific coefficient in the differential equation for a four-loop propagator over the first characteristic after removing an overall monomial factor.
+The FiniteFlow entry does not include a few hundred probes used for degree determinations.
+For the algorithm presented in section 3, the scaling powers in equation (10) are completely determined by the numerator in the present example.
+Thus, removing factors from the denominator does not affect the number of probes needed.
+However, we do observe a slight reduction in the number of required evaluations with FireFly, reducing the overhead to $26\%$ compared to the optimum.
+The number of evaluations needed with FiniteFlow exceeds the number of non-vanishing coefficients to be determined by about $43\%$.
 4.2 Diphoton Plus Jet Production at Two Loops
@@
 Specifically, we choose the parity-even contribution with a left-handed quark and a gluon in the initial state, a negative gluon helicity, opposite-sign photon helicities, and no closed fermion loops.
-Denoting the number of colours by $N_{C}$, we select the largest rational coefficient by Mathematica’s ByteCount in the part proportional to $N_{C}^{0}$.
+Denoting the quark helicity by $\lambda_{q}$, the number of active flavours by $n_{f}$, the number of colours by $N_{C}$, and the parity transformation operator by $\hat{P}$, the reduction has the structure
+[…]
+where $c_{-2},c_{0},c_{2}$ are linear combinations of pentagon functions [41] with rational coefficients.
+From $c_{0}$ we select the largest of these coefficients by Mathematica’s ByteCount.
 We write this coefficient as a rational function in $x_{23},x_{34},x_{45},x_{51}$, where $x_{ij}=\frac{s_{ij}}{s_{12}}$, and
 $s_{12}=(p_{1}+p_{2})^{2},s_{23}=(p_{2}-p_{3})^{2},s_{34}=(p_{3}+p_{4})^{2},s_{45}=(p_{4}+p_{5})^{2},s_{51}=(p_{1}-p_{5})^{2}$
-are Mandelstam invariants. In this example, the full coefficient can be reconstructed using a single prime field. The number of required probes is shown in table 2.
-Algorithm # Probes
-Homogeneous 248 688248688248\,688248 688
-Scaling 169 132169132169\,132169 132
-Optimal 30 4903049030\,49030 490
+are Mandelstam invariants.
+After determining the numerator and denominator degrees our ansatz according to equation (8) contains 136 934136\,934 unknown coefficients.
+However, the actual rational function is much sparser than in the example in section 4.1 and only approximately $22\%$ of these coefficients are non-zero.
+In this example, the full coefficient can be reconstructed using a single prime field.
+We collect the number of required probes in table 3.
+Method Number of Probes
+This work   169 132169\,132
+FireFly   163 094163\,094
+Optimal   30 49030\,490
 Table 3: Number of probes required to reconstruct the coefficient in the reduction of the two-loop diphoton plus jet amplitude.
-The scaling algorithm introduced in section 3 performs notably better than the homogeneous one, requiring approximately $32\%$ fewer probes.
-Still, both algorithms are far from optimal for this scenario and algorithms for sparse reconstruction [26, 27, 36, 37, 38] will likely perform much better.
+The scaling algorithm introduced in section 3 performs slightly worse than FireFly’s reconstruction.
+Both algorithms are far from optimal for this scenario, requiring more than five probes for each unknown coefficient.
+The number of required reconstruction probes after removing an overall monomial factor $x_{23}^{2}x_{34}^{2}x_{45}^{2}x_{51}^{2}$ is shown in table 4.
+As in section 4.1, we see no improvement for the implementation of the algorithm presented in this work.
+In contrast, FireFly needs approximately $20\%$ fewer function evaluations than before.
+Most strikingly, FiniteFlow is much closer to optimal than both FireFly and the scaling reconstruction implementation, especially when using the FFPolyVandermonde alternative polynomial reconstruction method.
+Even when enabling its identification of univariate factors, FireFly still requires 87 48587\,485 probes, substantially more than FiniteFlow.
+This difference between FiniteFlow and FireFly is unexpected and deserves closer inspection.
+However, since the focus of the present work is on the scaling algorithm and dense reconstruction, we leave further investigation to future work.
+Method Number of Probes
+This work   169 132169\,132
+FireFly   129 894129\,894
+FiniteFlow $\gtrsim$49\,216$\$
+FiniteFlow with FFPolyVandermonde $\gtrsim$47\,381$\$
+Optimal   30 49030\,490
+Table 4: Number of required probes after removing the overall monomial prefactor. The FiniteFlow entries do not include a few hundred probes used for degree determinations.
 5 Conclusion
 I have presented an algorithm for the reconstruction of dense multivariate rational functions.
-Multiple variables are mapped onto a single variable, using scaling powers and shifts chosen such that the mapping can be reversed.
+Multiple variables are mapped onto a single variable, using scaling powers and shifts chosen such that the mapping can be inverted.
 In this way, the problem is reduced to well-known univariate rational reconstruction.
-The algorithm is tested on two examples taken from complex reductions to basis integrals, a massive four-loop propagator and and two-loop five-point amplitude.
-For the dense rational function encountered in the four-loop problem, the required number of probes exceeds the number of unknown coefficients by only about $6\%$.
-In the sparse two-loop example, the overhead is much larger and the number of probes exceeds the number of non-vanishing coefficients by a factor of more than $5$.
-Still, in both cases the scaling algorithm compares favourably to dense homogeneous reconstruction [31].
-A promising avenue for future research would be to combine the univariate mapping with sparse rational reconstruction in a single variable [39].
+The algorithm is tested on two examples taken from complex reductions to basis integrals, a massive four-loop propagator and a two-loop five-point amplitude.
+For the dense rational function encountered in the four-loop problem, the required number of probes exceeds the number of unknown coefficients by only about $7\%$.
+This compares favourably with the current state-of-the-art programs FireFly [27, 29] and FiniteFlow [28].
+In the sparse two-loop example, the number of probes needed is about $4\%$ above the FireFly result when disabling factorisation.
+However, a comparison to FiniteFlow reveals that in this case there is substantial room for improvements for both FireFly and the presented algorithm.
+A further promising avenue for future research would be to combine the univariate mapping with sparse rational reconstruction in a single variable, see e.g.
+[42].
 Acknowledgements
-I thank P. Marquard for enlightening discussions and helpful comments on the manuscript.
+I thank P.
+Marquard for enlightening discussions and helpful comments on the manuscript.
+I further thank F.
+Lange and T.
+Peraro for communication on FireFly and FiniteFlow and for contributing code to the comparison in section 4.